1. A committee of 5 is to be formed from a group of 6 men and 5 women. if this individuals have an equal chance of being selected, what is the probability that the numbers are:
a. all men
b. two are men and three are women

Answer by srinivas.g(540)

total no of ways in which we can select a committee of 5 members from the group of 6 men and 5 women = 11 c 5
= 462
no of ways in which we can select a committee of 5 members from the group of 6 men = 6 c 5 = 6
the probability that all the members are men = 6/642
= 1/107


total no of ways in which we can select a committee of 2 men + 3 women from the group of 6 men and 5 women = 6c2 * 5c3
= 15 * 10
=150
he probability that the committee consist of 2 men & 3 women = 150/642
= 25/107